Circular No. 7397

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000bh IN ESO 573-14 J. Maza, University of Chile; and M. Hamuy, University of Arizona,

report the discovery by R. Antezana, on a 30-min unfiltered T-Max 400 film taken by L. González on Apr. 5.23 UT with the University of Chile’s Maksutov telescope, of a supernova (m_{pv} ∼ 16.0) located at α = 12^{h}21^{m}15^{s}.14, δ = −21^{o}59′55′′.7 (equinox 2000.0), which is ∼ 8′′ west and 11′′ south of the nucleus of the host galaxy (ESO 573-14, z = 0.023). The supernova was confirmed by M. M. Phillips (Carnegie Institution of Washington) from J and H images taken with the Las Campanas 1.02-m Swope telescope on Apr. 6, and by T. B. Williams (Rutgers University) from BVRI images taken with the Cerro Tololo Interamerican Observatory 0.9-m telescope on Apr. 7.22. Using the 2.54-m DuPont telescope at Las Campanas Observatory on Apr. 7, L. C. Ho (Carnegie Institution of Washington) obtained a spectrum (range 320–900 nm) that reveals SN 2000bh to be a type-Ia event near maximum light.

XTE J1118+480 L. Cook, Concord, CA; J. Patterson, Columbia University; D. Buczyn-

ski, Condor Brow Observatory; and R. Fried, Braeside Observatory, report for the Center for Backyard Astrophysics (cf. Patterson 1998, Sky Tel. 96(4), 77) that a 5-night V -band light curve of the optical counterpart (IAUC 7390) shows a regular periodic variation with a period of 0.1706(9) day and a full amplitude of 0.061(5) mag. The signal is smooth and sinusoidal, suggesting an x-ray heating effect in a close binary with an orbital period of 0.1706 day.

Visual magnitude estimates: Mar. 31.844 UT, 13.0 (P. Schmeer, Bischmisheim, Germany); Apr. 2.123, 12.9 (Schmeer); 2.902, 12.9 (Schmeer); 4.010, 13.0 (Schmeer); 6.838, 12.7 (K. Hornoch, Lelekovice, Czech Rep.).

SUPERNOVA 2000E IN NGC 6951 CCD magnitudes, unfiltered unless otherwise noted: Jan. 28.740 UT,

R = 14.2 ± 0.1 (L. Kiss, K. Sárneczky, and K. Sziládi, Piszkéstető station, Konkoly Observatory); 28.816, 14.3 (S. Yoshida and K. Kadota, Ageo, Saitama, Japan); Mar. 16.11, 16.0 ± 0.1 (Sárneczky, Sziládi, and Kiss); 29.024, 15.8 (Hornoch); Apr. 3.028, 16.0 (Hornoch).

2000 April 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green